# Diversified and Adaptive Negative Sampling on Knowledge Graphs

### v1

@@ Abstract
keywords:
Knowledge graphs, Graph representation learning, Graph neural networks, Negative sampling
journal: Neural Networks
[…]
{highlights}
We design a two-way generator to produce diverse negative triplets, to increase the overall informativeness.
We employ a FiLM layer to adapt the global generator model into local models, to increase the individual informativeness of the negative triplets.
We conduct extensive experiments on three benchmark datasets. The results demonstrate the superiority of our proposed approach.
@@ 1 Introduction
In this paper, we focus on the problem of knowledge graph embedding. The high-level idea is that the embedding vectors of entities and relations co-occurring in the same fact should be bounded by certain constraints due to their relatedness. For instance, consider a fact $\tau=\langle h=\mathtt{Beijing},r=\mathtt{isCapitalOf},t=\mathtt{China}\rangle$ and a classic method TransE [2]. TransE maps each entity and relation in the fact to vectors, i.e., $\mathbf{e}_{h}$, $\mathbf{e}_{r}$, $\mathbf{e}_{t}$, respectively, so that they approximately satisfy the constraint $\mathbf{e}_{h}+\mathbf{e}_{r}\approx\mathbf{e}_{t}$ by minimizing the loss $\|\mathbf{e}_{h}+\mathbf{e}_{r}-\mathbf{e}_{t}\|$. On the contrary, a nonfact such as $\langle h=\mathtt{Beijing},r=\mathtt{isCapitalOf},t^{\prime}=\mathtt{Russia}\rangle$ would maximize the loss $\|\mathbf{e}_{h}+\mathbf{e}_{r}-\mathbf{e}_{t^{\prime}}\|$. Given this contrast, the factual triplets are known as positive triplets (or examples), whereas the non-factual triplets are called negative triplets. Although positive triplets are readily available, negative triplets are often obtained through random sampling. More recent works [43, 3, 1, 48, 30, 50] explore advanced constraints or losses [2, 36, 44] on the triplets, but the sampling strategy for negative triplets remains a crucial yet less explored problem.
@@
In view of the above, we propose a Diversified and Adaptive Negative Sampling (DANS) approach for knowledge graph embedding, to improve both the overall and individual informativeness of negative triplets. Similar to previous state-of-the-art approaches such as KBGAN [3], we adopt a generative adversarial network (GAN) [38] for the generation of negative samples. However, there are two significant differences from previous GAN-based negative sampling on the knowledge graph. First, we design a two-way generator to produce diversified samples that are associated with both entities and relations w.r.t. a positive triplet, which aims to increase the overall informativeness of the samples. More specifically, the generator consists of two pathways to produce two different kinds of negative triplets associated with a given entity and entity-relation, respectively. Second, we design an adaptive mechanism to modulate the global generator model into local models to handle the differences across entities and relations, which aims to increase the individual informativeness of the samples in a finer-grained manner. In particular, we employ a Feature-wise Linear Modulation (FiLM) layer [26] that conditions the generator on a given entity or entity-relation input. In summary, we make the following contributions:
  • We design a two-way generator to produce diverse negative triplets, to increase the overall informativeness.
  • We employ a FiLM layer to adapt the global generator model into local models, to increase the individual informativeness of the negative triplets.
  • We conduct extensive experiments on three benchmark datasets. The results demonstrate the superiority of our proposed approach.
@@ 2 Background
Negative sampling is an important issue in various machine learning tasks such as recommendation systems [28] and natural language processing [20]. In the context of knowledge graph embedding, negative triplets are often constructed by replacing the tail or head entity in a positive triplet with a randomly sampled entity [2, 40, 17]. Unfortunately, in uniform [2] or popularity-weighted sampling [20], the sampled entity could be completely unrelated to the head or the relation, and therefore be less informative.
To sample more informative negative triplets, researchers have leveraged different heuristics or learning strategies. Several structure-aware models [1, 46, 18] exploit the graph structures, which generally select negative examples in the neighborhood of positive examples. For example, SANS [1] hypothesizes that entities that are in close proximity to each other, but do not share a direct relationship, are better candidates for negative sampling. In a similar spirit, PinSage [18] generates localized graphs via random walks to extract informative negative samples. However, these approaches have a high risk of selecting false negatives, as not explicitly related entities in close proximity could still form positive triplets due to the incompleteness of the observed graph.
Other approaches seek to quantify the informativeness of the negative triplets through various learning strategies, including GANs [3, 38, 11, 47], reinforcement learning [39, 49], and importance sampling [48]. These methods provide a more explicit and systematic scoring of negative triplets which often led to better performance. However, these approaches do not consider the diversity and adaptiveness of negative sampling, which are crucial to the overall and individual informativeness of the negative triplets, respectively.
@@ 3.1 Problem formulation and preliminaries
where $y_{\tau}=1$ if $\tau\in\mathcal{E}$, else $y_{\tau}=0$. We implement $\mathcal{F}$ using three popular decoders, namely, DistMult [44], ComplEx [36] and RotatE [34]. We provide the DistMult function below, and leave the details of ComplEx and RotatE to A.
@@ 3.2 Adaptive two-way generator
To sum up, the trainable parameters in the adaptive two-way generator, $\Theta_{G}$, include the weights in the two global pathways and the FiLM layer weights for each layer in each pathway. Assuming a total of $M$ hidden layers in the global pathways, we would have $\Theta_{G}=\{\Theta_{G_{E}},\Theta_{G_{R}}\}\cup\{\Theta_{F_{E},\alpha}^{m},\Theta_{F_{E},\beta}^{m},\Theta_{F_{R},\alpha}^{m},\Theta_{F_{R},\beta}^{m}:m\leq M\}$.
@@ 3.3 Two-way discriminator
As in a standard GAN architecture, a discriminator is needed to help the generator produce high-quality fake entities that mimic real entities. Specifically, the discriminator and the generator compete with each other in a minimax game, in which the generator aims to fool the discriminator by producing realistic looking entities, while the discriminator aims to beat the generator by distinguishing the real and fake entities. In our case, given the two-way generator, we further equip the discriminator with the ability to distinguish the fake entities generated by the two pathways, which can further differentiate and diversify the two pathways.
@@
[…]
@@
$\mathcal{L}_{\text{Adv}}(\hat{y}_{\text{Adv},i},y_{\text{Adv},i})=-y_{\text{Adv},i}\log\hat{y}_{\text{Adv},i}$
$-(1-y_{\text{Adv},i})\log(1-\hat{y}_{\text{Adv},i}),$ (11)
$\mathcal{L}_{\text{Aux}}(\hat{y}_{\text{Aux},i},y_{\text{Aux},i})=-y_{\text{Aux},i}\log\hat{y}_{\text{Aux},i}$
$-(1-y_{\text{Aux},i})\log(1-\hat{y}_{\text{Aux},i}),$ (12)
@@ 3.4 Adversarial training
$\textstyle\arg\max_{\Theta_{G}}$ $\ \mathbb{E}_{t}{\mathcal{L}_{\text{Adv}}(\hat{y}_{\text{Adv},t},1)}$
$+\mathbb{E}_{t^{\prime}}{\mathcal{L}_{\text{Adv}}(\hat{y}_{\text{Adv},t^{\prime}},0)}+\mathbb{E}_{t^{\prime\prime}}{\mathcal{L}_{\text{Adv}}(\hat{y}_{\text{Adv},t^{\prime\prime}},0)}$
@@
where $t$ is a real tail entity, and $t^{\prime},t^{\prime\prime}$ are fake tail entities from $G_{E}$ and $G_{R}$, respectively (again, we only illustrate the case where the tail entity in a positive triplet is replaced). The last term in Eq. (3.4) is a regularization term on the scaling and shifting factors to prevent overfitting as in Oreshkin et al. [24], and $\lambda$ is a hyper-parameter to control the strength of regularization. On the other hand, the goal of the discriminator is to overcome the generators by distinguishing fake and real entities, as well as fake entities from different generator pathways, as follows.
$\textstyle\arg\min_{\Theta_{D}}$ $\ \mathbb{E}_{i}{\mathcal{L}_{\text{Adv}}(\hat{y}_{\text{Adv},i},y_{\text{Adv},i})}$
$+$ $\ \mathbb{E}_{i\neq t}{\mathcal{L}_{\text{Aux}}(\hat{y}_{\text{Aux},i},y_{\text{Aux},i})},$ (14)
@@
Following a typical adversarial training scheme in negative sampling on knowledge graphs in KBGAN [3], we alternate the model updating among the three parties, as follows. First, we train the generator by updating the generator parameters $\Theta_{G}$ with Eq. (3.4), while freezing the discriminator parameters $\Theta_{D}$ and the base model parameters $\Theta$. Next, we update $\Theta_{D}$ with Eq. (3.4), while freezing $\Theta_{G},\Theta$. Finally, we update $\Theta$ by minimizing the loss on the positive and negative triples in Eq. (2), while freezing the other two parameter sets. We repeat the three steps until the convergence of all parties are achieved.
@@ 4 Experiments
We perform empirical evaluation on three benchmark knowledge graphs. We first compare the empirical performance of the proposed model DANS [^1] We include the code for review in Supplementary Materials. with state-of-the-art baselines. In addition, we seek to address a number of research questions (RQ) through more in-depth empirical analysis. RQ1: Does the two-way design in the generator improve model performance? RQ2: Does the adaptive FiLM layer in the generator improve model performance? RQ3: What is the impact of the number of negative triplets and adaptive regularization, respectively? RQ4: Can we observe the diversity and adaptiveness of generated triplets?
@@ 4.1 Experimental Design
Entities Relations Train Val Test Total
@@
Datasets. Three benchmark knowledge graphs are used for our experiment. (1) WN18RR [2], a harder variant of WN18 [7], which is derived from WordNet consisting of hyponym and hypernym relations between words. Compared to WN18, WN18RR removes inverse relations to minimize leakage from training. (2) NELL-995 [43] is a subset of the web-based facts collected by the 995th iteration of the Nell system [5] which contains a large pool of entity types and only the top 200 relations are retained. (3) UMLS [7] is a specialized knowledge base containing medical entities and their semantic relationships. The entities are biomedical concepts (e.g., disease, antibiotic), and the relations include interactions such as $\mathtt{treats}$ and $\mathtt{diagnosis}$. Table 1 gives a summary of the datasets used.
Task and evaluation. We employ the standard knowledge graph completion task [21, 4, 13, 6]. Specifically, for each positive test triplet, we construct a list of candidate triplets that also include negative triplets, which are obtained by replacing either the head or tail of the positive triplet with every other entity in the dataset. To avoid false negatives, we follow previous work Bordes et al. [2] by adopting their “filtered setting”. We then rank the candidate triplets based on the scoring function. For evaluation, we adopt several standard ranking metrics including Mean Reciprocal Ranking (MRR), Hit ratio at 1 (H@1) and Normalized discounted cumulative gain at 5 (NDCG@5) [35]. Details of these ranking metrics can be found in B.
@@
(1) Negative samplers with the same RGCN backbone [29] and decoders. In other words, they are flexible “plug-ins” that only replace the sampling strategy for a fair comparison to our method DANS. They include Rand, which replaces the head or tail entity with a uniformly sampled random entity; Pop [20]: a variant of Rand that substitutes uniform sampling with popularity-weighted sampling; Self-adv [35]: a self-adversarial negative sampling methodology; MCNS [45]: a model which derives negative samples from a distribution that is positively but sub-linearly correlated with the positive distribution.
@@
In addition, the hyper-parameters related to negative sampling via Metr-opolis-Hastings in RGCN-MCNS have been copied from the original paper ’s link prediction experiments in Yang et al. [45]. To reduce the variance resulting from parameter initialization, the experimental results are calculated from an average of five runs with different seeds in all methods. Furthermore, every method is standardized to use the triplet loss in Eqs.(2). Other baseline settings have also been tuned according to the recommendations of the literature. Additional details can be found in C.
@@ 4.2 Results and Analysis
{tabu}
@c—ccccccccc@ Sampling WN18RR NELL-995 UMLS
@@
Rand .372$\pm$.002 .343$\pm$.003 .369$\pm$.005 .218$\pm$.001 .146$\pm$.002 .219$\pm$.002 .696$\pm$.010 .607$\pm$.082 .693$\pm$.007
Pop .374$\pm$.002 .342$\pm$.002 .376$\pm$.006 .216$\pm$.001 .142$\pm$.002 .216$\pm$.003 .680$\pm$.009 .589$\pm$.012 .692$\pm$.005
Self-adv .370$\pm$.007 .332$\pm$.010 .373$\pm$.006 .238$\pm$.003 .156$\pm$.003 .241$\pm$.003 .717$\pm$.009 .624$\pm$.015 .733$\pm$.008
MCNS .376$\pm$.004 .340$\pm$.005 .374$\pm$.006 .226$\pm$.002 .144$\pm$.002 .221$\pm$.003 .700$\pm$.002 .606$\pm$.008 .717$\pm$.002
DANS .381$\pm$.006 .352$\pm$.007 .386$\pm$.008 .227$\pm$.004 .162$\pm$.007 .220$\pm$.009 .724$\pm$.008 .641$\pm$.009 .725$\pm$.008
@@
Rand .234$\pm$.009 .110$\pm$.003 .260$\pm$.008 .182$\pm$.003 .093$\pm$.003 1̇89$\pm$.003 .817$\pm$.015 .683$\pm$.021 .855$\pm$.013
Pop .235$\pm$.007 .095$\pm$.003 .268$\pm$.007 .181$\pm$.002 .131$\pm$.002 .200$\pm$.003 .800$\pm$.005 .673$\pm$.010 .839$\pm$.004
Self-adv .202$\pm$.007 .058$\pm$.010 .235$\pm$.006 .186$\pm$.002 .096$\pm$.003 .194$\pm$.002 .809$\pm$.007 .677$\pm$.007 .848$\pm$.007
MCNS .242$\pm$.009 .132$\pm$.004 .288$\pm$.006 .194$\pm$.003 .122$\pm$.004 .200$\pm$.004 .822$\pm$.005 .682$\pm$.006 .884$\pm$.006
DANS .249$\pm$.002 .154$\pm$.001 .274$\pm$.003 .195$\pm$.010 .135$\pm$.011 .208$\pm$.010 .833$\pm$.004 .716$\pm$.006 .866$\pm$.005
@@
Rand .386$\pm$.007 .346$\pm$.005 .390$\pm$.006 .245$\pm$.004 .172$\pm$.003 2̇51$\pm$.006 .898$\pm$.008 .822$\pm$.017 .920$\pm$.015
Pop .389$\pm$.011 .341$\pm$.007 .387.$\pm$.012 .241$\pm$.005 .179$\pm$.006 .245$\pm$.004 .840$\pm$.009 .747$\pm$.009 .865$\pm$.008
Self-adv .375$\pm$.006 .329$\pm$.011 .382$\pm$.013 .250$\pm$.005 .181$\pm$.007 .277$\pm$.008 .908$\pm$.009 .844$\pm$.006 .925$\pm$.010
MCNS .392$\pm$.008 .343$\pm$.007 .394$\pm$.008 .248$\pm$.007 .177$\pm$.004 .264$\pm$.009 .879$\pm$.007 .835$\pm$.005 .892$\pm$.011
DANS .404$\pm$.005 .347$\pm$.004 .392$\pm$.009 .257$\pm$.006 .186$\pm$.010 .255$\pm$.008 .920$\pm$.007 .857$\pm$.011 .927$\pm$.008
Table 3: Performance comparison with baselines (all using the DistMult decoder). See Table 4.2 caption for entry styles.
@@
SANS-RW .349$\pm$.010 .340$\pm$.013 .334$\pm$.010 .135$\pm$.006 .109$\pm$.008 .110$\pm$.008 .510$\pm$.008 .369$\pm$.009 .478$\pm$.003
NSCaching .374$\pm$.002 .337$\pm$.003 .374$\pm$.002 .177$\pm$.004 .150$\pm$.003 .140$\pm$.002 .625$\pm$.004 .508$\pm$.021 .607$\pm$.004
KBGAN .172$\pm$.004 .070$\pm$.006 .155$\pm$.002 .170$\pm$.002 .077$\pm$.004 .195$\pm$.009 .680$\pm$.005 .556$\pm$.023 .654$\pm$.004
CAKE .353$\pm$.007 .345$\pm$.005 .351$\pm$.008 .204$\pm$.006 .130$\pm$.007 .175$\pm$.012 .441$\pm$.013 .365$\pm$.008 .383$\pm$.010
SMiLE .315$\pm$.006 .291$\pm$.007 .294$\pm$.012 .131$\pm$.004 .127$\pm$.005 .105$\pm$.008 .414$\pm$.015 .345$\pm$.007 .372$\pm$.013
DANS .381$\pm$.006 .352$\pm$.007 .386$\pm$.008 .227$\pm$.004 .162$\pm$.007 .220$\pm$.009 .724$\pm$.008 .641$\pm$.009 .725$\pm$.008
Table 4.2 reports quantitative comparison against the first category of baselines involving different negative samplers under the same backbone and decoder. Overall, our model DANS consistently leads to better performance for DistMult, RotatE and ComplEx decoders. This shows the robustness of our approach across various decoders. In general, DANS performs better than Rand and its variant Pop, showing that it is important to account for the informativeness of negative triples which are missing in random and popularity-weighted sampling. Since Self-Adv accounts for the informativeness by giving more weight to higher quality triplets, it generally outperforms Rand and Pop. It still lags behind DANS in most cases as it ignores the concepts of diversity and adaptiveness. The variant MCNS shows better performance than Rand and its variant Pop but loses to DANS  as MCNS was originally designed for homogeneous graphs.
Overall, DANS has obtained favourable performance, showing the importance of diversity and adaptiveness during negative sampling. We will conduct further ablation study in the next part to examine the contribution from each aspect. Finally, we have included the experimental results for dataset FB15k-237 which show favorable performance on ComplEx decoder in D.

### v2

@@ 1 Introduction
In this paper, we focus on the problem of knowledge graph embedding. The high-level idea is that the embedding vectors of entities and relations co-occurring in the same fact should be bounded by certain constraints due to their relatedness. For instance, consider a fact $\tau=\langle h=\mathtt{Beijing},r=\mathtt{isCapitalOf},t=\mathtt{China}\rangle$ and a classic method TransE [2]. TransE maps each entity and relation in the fact to vectors, i.e., $\mathbf{e}_{h}$, $\mathbf{e}_{r}$, $\mathbf{e}_{t}$, respectively, so that they approximately satisfy the constraint $\mathbf{e}_{h}+\mathbf{e}_{r}\approx\mathbf{e}_{t}$ by minimizing the loss $\|\mathbf{e}_{h}+\mathbf{e}_{r}-\mathbf{e}_{t}\|$. On the contrary, a nonfact such as $\langle h=\mathtt{Beijing},r=\mathtt{isCapitalOf},t^{\prime}=\mathtt{Russia}\rangle$ would maximize the loss $\|\mathbf{e}_{h}+\mathbf{e}_{r}-\mathbf{e}_{t^{\prime}}\|$. Given this contrast, the factual triplets are known as positive triplets (or examples), whereas the non-factual triplets are called negative triplets. Although positive triplets are readily available, negative triplets are often obtained through random sampling. More recent works [34, 3] explore advanced constraints or losses [2, 29] on the triplets, but the sampling strategy for negative triplets remains a crucial yet less explored problem.
@@
In view of the above, we propose a Diversified and Adaptive Negative Sampling (DANS) approach for knowledge graph embedding, to improve both the overall and individual informativeness of negative triplets. Similar to previous state-of-the-art approaches such as KBGAN [3], we adopt a generative adversarial network (GAN) [30] for the generation of negative samples. However, there are two significant differences from previous GAN-based negative sampling on the knowledge graph. First, we design a two-way generator to produce diversified samples that are associated with both entities and relations w.r.t. a positive triplet, which aims to increase the overall informativeness of the samples. More specifically, the generator consists of two pathways to produce two different kinds of negative triplets associated with a given entity and entity-relation, respectively. Second, we design an adaptive mechanism to modulate the global generator model into local models to handle the differences across entities and relations, which aims to increase the individual informativeness of the samples in a finer-grained manner. In particular, we employ a Feature-wise Linear Modulation (FiLM) layer [23] that conditions the generator on a given entity or entity-relation input.
In summary, we make the following contributions. (1) We design a two-way generator to produce diverse negative triplets, to increase the overall informativeness. (2) We employ a FiLM layer to adapt the global generator model into local models, to increase the individual informativeness of the negative triplets. (3) We conduct extensive experiments on three benchmark datasets. The results demonstrate the superiority of our proposed approach.
@@ 2 Background
Negative sampling is an important issue in various machine learning tasks such as recommendation systems [25] and natural language processing [17]. In the context of knowledge graph embedding, negative triplets are often constructed by replacing the tail or head entity in a positive triplet with a randomly sampled entity [2]. Unfortunately, in uniform [2] or popularity-weighted sampling [17], the sampled entity could be completely unrelated to the head or the relation, and therefore be less informative.
To sample more informative negative triplets, researchers have leveraged different heuristics or learning strategies. Structure-aware models [1, 37] exploit the graph structures, which generally select negative examples in the neighborhood of positive examples. For example, SANS [1] hypothesizes that entities that are in close proximity to each other, but do not share a direct relationship, are better candidates for negative sampling. In a similar spirit, PinSage [14] generates localized graphs via random walks to extract informative negative samples. However, these approaches have a high risk of selecting false negatives, as not explicitly related entities in close proximity could still form positive triplets due to the incompleteness of the observed graph.
Other approaches seek to quantify the informativeness of the negative triplets through various learning strategies, including GANs [3, 30], reinforcement learning [31], and importance sampling [39]. These methods provide a more explicit and systematic scoring of negative triplets which often led to better performance. However, these approaches do not consider the diversity and adaptiveness of negative sampling, which are crucial to the overall and individual informativeness of the negative triplets, respectively.
@@ 3.1 Problem formulation and preliminaries
where $y_{\tau}=1$ if $\tau\in\mathcal{E}$, else $y_{\tau}=0$. We implement $\mathcal{F}$ using three popular decoders, namely, DistMult [35], ComplEx [29] and RotatE [27]. We provide the DistMult function below, and leave the details of ComplEx and RotatE to Appendix A of the extended version [15].
@@ 3.2 Adaptive two-way generator
To sum up, the trainable parameters in the adaptive two-way generator, $\Theta_{G}$, include the weights in the two global pathways and the FiLM layer weights for each layer in each pathway. For a total of $M$ hidden layers in the global pathways, we have $\Theta_{G}=\{\Theta_{G_{E}},\Theta_{G_{R}}\}\cup\{\Theta_{F_{E},\alpha}^{m},\Theta_{F_{E},\beta}^{m},\Theta_{F_{R},\alpha}^{m},\Theta_{F_{R},\beta}^{m}:m\leq M\}$.
@@ 3.3 Two-way discriminator
As in a standard GAN, a discriminator is needed to help the generator produce high-quality fake entities that mimic real entities. Specifically, the discriminator and the generator compete with each other in a minimax game, in which the generator aims to fool the discriminator by producing realistic looking entities, while the discriminator aims to beat the generator by distinguishing the real and fake entities. In our case, given the two-way generator, we further equip the discriminator with the ability to distinguish the fake entities generated by the two pathways, which can further differentiate and diversify the two pathways.
@@
where $D_{\text{Adv}}$ and $D_{\text{Aux}}$ are implemented as a fully connected layer, and $\tilde{\mathbf{e}}_{i}=\textsc{{Mlp}}(\mathbf{e}_{i};\Theta_{D_{S}})$ is a shared hidden representation computed from the embedding $\mathbf{e}_{i}$ of a real or fake entity $i$. The shared hidden representation allows both $D_{\text{Adv}}$ and $D_{\text{Aux}}$ to benefit from each other during training [20], as they collectively try to distinguish three different classes of samples ($t$, $^{\prime}, t′′t^{\prime\prime}$).
@@
$\mathcal{L}_{\text{Adv}}(\hat{y}_{\text{Adv},i},y_{\text{Adv},i})=-y_{\text{Adv},i}\log\hat{y}_{\text{Adv},i}-(1-y_{\text{Adv},i})\log(1-\hat{y}_{\text{Adv},i}),$ (11)
$\mathcal{L}_{\text{Aux}}(\hat{y}_{\text{Aux},i},y_{\text{Aux},i})=-y_{\text{Aux},i}\log\hat{y}_{\text{Aux},i}-(1-y_{\text{Aux},i})\log(1-\hat{y}_{\text{Aux},i}),$ (12)
@@ 3.4 Adversarial training
$\textstyle\arg\max_{\Theta_{G}}$ $\ \mathbb{E}_{t}{\mathcal{L}_{\text{Adv}}(\hat{y}_{\text{Adv},t},1)}+\mathbb{E}_{t^{\prime}}{\mathcal{L}_{\text{Adv}}(\hat{y}_{\text{Adv},t^{\prime}},0)}+\mathbb{E}_{t^{\prime\prime}}{\mathcal{L}_{\text{Adv}}(\hat{y}_{\text{Adv},t^{\prime\prime}},0)}$
@@
where $t$ is a real tail entity, and $t^{\prime},t^{\prime\prime}$ are fake tail entities from $G_{E}$ and $G_{R}$, respectively (again, we only illustrate the case where the tail entity in a positive triplet is replaced). The last term in Eq. (3.4) is a regularization term on the scaling and shifting factors to prevent overfitting [21], and $\lambda$ is a hyper-parameter to control the strength of regularization. On the other hand, the goal of the discriminator is to overcome the generators by distinguishing fake and real entities, as well as fake entities from different generator pathways, as follows.
$\textstyle\arg\min_{\Theta_{D}}$ $\ \mathbb{E}_{i}{\mathcal{L}_{\text{Adv}}(\hat{y}_{\text{Adv},i},y_{\text{Adv},i})}+\mathbb{E}_{i\neq t}{\mathcal{L}_{\text{Aux}}(\hat{y}_{\text{Aux},i},y_{\text{Aux},i})},$ (14)
@@
Following a typical adversarial training scheme in negative sampling on knowledge graphs in KBGAN [3], we alternate the model updating among the three parties, as follows. First, we train the generator by updating the generator parameters $\Theta_{G}$ with Eq. (3.4), while freezing the discriminator parameters $\Theta_{D}$ and the base model parameters $\Theta$. Next, we update $\Theta_{D}$ with Eq. (14), while freezing $\Theta_{G},\Theta$. Finally, we update $\Theta$ by minimizing the loss on the positive and negative triples in Eq. (2), while freezing the other two parameter sets. We repeat the three steps until the convergence of all parties are achieved.
@@ 4 Experiments
We perform empirical evaluation on three benchmark knowledge graphs. We first compare the empirical performance of the proposed model DANS [^1] https://github.com/liuran998/DANS with state-of-the-art baselines. In addition, we seek to address a number of research questions (RQs). RQ1: Does the two-way design in the generator improve model performance? RQ2: Does the adaptive FiLM layer in the generator improve model performance? RQ3: What is the impact of the number of negative triplets and adaptive regularization, respectively? RQ4: Can we observe the diversity of generated triplets?
@@ 4.1 Experimental Design
Entities Relations Train Validation Test Total
@@
Datasets. Three benchmark knowledge graphs are used for our experiment. (1) WN18RR [2], a harder variant of WN18 [7], which is derived from WordNet consisting of hyponym and hypernym relations between words. Compared to WN18, WN18RR removes inverse relations to minimize leakage from training. (2) NELL-995 [8] is a subset of the web-based facts collected by the 995th iteration of the Nell system [5] which contains a large pool of entity types and only the top 200 relations are retained. (3) UMLS [7] is a specialized knowledge base containing medical entities and their semantic relationships. The entities are biomedical concepts (e.g., disease, antibiotic), and the relations include interactions such as $\mathtt{treats}$ and $\mathtt{diagnosis}$. Table 1 gives a summary of the datasets.
Task and evaluation. We employ the standard knowledge graph completion task [18, 4, 12, 6]. Specifically, for each positive test triplet, we construct a list of candidate triplets that also include negative triplets, which are obtained by replacing either the head or tail of the positive triplet with every other entity in the dataset. To avoid false negatives, we follow previous work [2] by adopting their “filtered setting”. We then rank the candidate triplets based on the scoring function. For evaluation, we adopt several standard ranking metrics including Mean Reciprocal Ranking (MRR), Hit ratio at 1 (H@1) and Normalized discounted cumulative gain at 5 (NDCG@5) [28]. Details of these ranking metrics can be found in Appendix B of the extended version [15].
@@
(1) Negative samplers with the same RGCN backbone [26] and decoders. In other words, they are flexible “plug-ins" that only replace the sampling strategy for a fair comparison to our method DANS. They include Rand, which replaces the head or tail entity with a uniformly sampled random entity; Pop [17]: a variant of Rand that substitutes uniform sampling with popularity-weighted sampling; Self-adv [28]: a self-adversarial negative sampling methodology; MCNS [36]: a model which derives negative samples from a distribution that is positively but sub-linearly correlated with the positive distribution.
@@
In addition, the hyper-parameters related to negative sampling via Metropolis-Hastings in RGCN-MCNS follow the original paper’s link prediction experiments [36]. To reduce the variance resulting from parameter initialization, the experimental results are calculated from an average of five runs with different seeds in all methods. Furthermore, every method is standardized to use the triplet loss in Eqs.(2). Other baseline settings have also been tuned according to the recommendations of the literature. Additional details can be found Appendix C of the extended version [15].
@@ 4.2 Results and Analysis
Table 2 reports quantitative comparison against the first category of baselines involving different negative samplers under the same backbone and decoder. Overall, our model DANS consistently leads to better performance for DistMult, RotatE and ComplEx decoders. This shows the robustness of our approach across various decoders. In general, DANS performs better than Rand and its variant Pop, showing that it is important to account for the informativeness of negative triples which are missing in random and popularity-weighted sampling. Since Self-Adv accounts for the informativeness by giving more weight to higher quality triplets, it generally outperforms Rand and Pop. It still lags behind DANS in most cases as it ignores the concepts of diversity and adaptiveness. The variant MCNS shows better performance than Rand and its variant Pop but loses to DANS  as MCNS was originally designed for homogeneous graphs.
@@
Sampling WN18RR NELL-995 UMLS
@@
Rand .372±.002 .343±.003 .369±.005 .218±.001 .146±.002 .219±.002 .696±.010 .607±.082 .693±.007
Pop .374±.002 .342±.002 .376±.006 .216±.001 .142±.002 .216±.003 .680±.009 .589±.012 .692±.005
Self-adv .370±.007 .332±.010 .373±.006 .238±.003 .156±.003 .241±.003 .717±.009 .624±.015 .733±.008
MCNS .376±.004 .340±.005 .374±.006 .226±.002 .144±.002 .221±.003 .700±.002 .606±.008 .717±.002
DANS .381±.006 .352±.007 .386±.008 .227±.004 .162±.007 .220±.009 .724±.008 .641±.009 .725±.008
@@
Rand .234±.009 .110±.003 .260±.008 .182±.003 .093±.003 1̇89±.003 .817±.015 .683±.021 .855±.013
Pop .235±.007 .095±.003 .268±.007 .181±.002 .131±.002 .200±.003 .800±.005 .673±.010 .839±.004
Self-adv .202±.007 .058±.010 .235±.006 .186±.002 .096±.003 .194±.002 .809±.007 .677±.007 .848±.007
MCNS .242±.009 .132±.004 .288±.006 .194±.003 .122±.004 .200±.004 .822±.005 .682±.006 .884±.006
DANS .249±.002 .154±.001 .274±.003 .195±.010 .135±.011 .208±.010 .833±.004 .716±.006 .866±.005
@@
Rand .386±.007 .346±.005 .390±.006 .245±.004 .172±.003 2̇51±.006 .898±.008 .822±.017 .920±.015
Pop .389±.011 .341±.007 .387±.012 .241±.005 .179±.006 .245±.004 .840±.009 .747±.009 .865±.008
Self-adv .375±.006 .329±.011 .382±.013 .250±.005 .181±.007 .277±.008 .908±.009 .844±.006 .925±.010
MCNS .392±.008 .343±.007 .394±.008 .248±.007 .177±.004 .264±.009 .879±.007 .835±.005 .892±.011
DANS .404±.005 .347±.004 .392±.009 .257±.006 .186±.010 .255±.008 .920±.007 .857±.011 .927±.008
Table 3: Performance comparison with baselines (all using the DistMult decoder). See Table 2 caption for entry styles.
@@
SANS-RW .349±.010 .340±.013 .334±.010 .135±.006 .109±.008 .110±.008 .510±.008 .369±.009 .478±.003
NSCaching .374±.002 .337±.003 .374±.002 .177±.004 .150±.003 .140±.002 .625±.004 .508±.021 .607±.004
KBGAN .172±.004 .070±.006 .155±.002 .170±.002 .077±.004 .195±.009 .680±.005 .556±.023 .654±.004
CAKE .353±.007 .345±.005 .351±.008 .204±.006 .130±.007 .175±.012 .441±.013 .365±.008 .383±.010
SMiLE .315±.006 .291±.007 .294±.012 .131±.004 .127±.005 .105±.008 .414±.015 .345±.007 .372±.013
DANS .381±.006 .352±.007 .386±.008 .227±.004 .162±.007 .220±.009 .724±.008 .641±.009 .725±.008
Overall, DANS has obtained favourable performance, showing the importance of diversity and adaptiveness during negative sampling. We will conduct further ablation study in the next part to examine the contribution from each aspect. Finally, we have included the experimental results for dataset FB15k-237 which show favorable performance on ComplEx decoder in Appendix D of the extended version [15].